OpenStudy (anonymous):

Help!! give medals :) Find the surface area of the triangular prism formed by the net (the volume is 3.46)

OpenStudy (anonymous):



OpenStudy (espex):

So you know your volume is \(\frac{1}{2}*b*h*l\), are the other dimensions you have in your drawing given to you?

OpenStudy (espex):

If so, all you need do is solve for your 'l' in this equation, this in turn will give you all the dimensions necessary to solve for your surface area.

OpenStudy (anonymous):

All the edges are 2 and then the triangle is give im still don't understand @eSpeX

OpenStudy (espex):

You have all the information you need to solve this. They have given you all of the edges, so you just need to plug the information in. What is your 'b', 'h', and 'l'?

OpenStudy (anonymous):

the bace and hight are 2 and would the lengs be 3sqrt

OpenStudy (espex):

Good, so now you can use those to find the area of the two triangles, add them to the area of the three rectangles, and you have your total surface area. The area of each triangle is what? The area of each rectangle is what?

OpenStudy (anonymous):

area of the 2 triangle is 13.86 are of the 3 rectangle is 12 then do I add them together??

OpenStudy (espex):

How did you get 13.86 for the area of the two triangles?

OpenStudy (anonymous):

i took 2X2X3sqrt =6.93X2=13.86

OpenStudy (espex):

The area of a triangle is \(\frac{1}{2}*b*h\) and you have two of them.

OpenStudy (anonymous):

so the area of one triangle would be 2 2x2=4 is the area of both triangles

OpenStudy (espex):

As you have an equilateral triangle, your area is going to be \(\frac{1}{2} * 2 * \sqrt{3}\) for each of your triangles.

OpenStudy (anonymous):

Ok thanks for the help i think i got it from here much apprecated

OpenStudy (espex):

You're welcome.